OpenStudy (anonymous):

Will medal A triangle has a perimeter of 18 inches. If one side has length 8 inches, find the length of the shortest side if the other lengths are in the ratio of 2 units : 3 units.

OpenStudy (anonymous):

a. 3 1/3 b. 4 c. 6 d. 7 1/5

OpenStudy (anonymous):

Let x be the shortest side. Solve the following for x:\[\frac{x}{10-x}=\frac{2}{3} \]\[x=4 \]